# HTMformer: Hybrid Time and Multivariate Transformer for Time Series Forecasting

### v2

@@ 4.1 Experiment Settings
All models are evaluated under identical settings, following the Time-Series-Library (Wang et al., 2024b), to ensure the fairness of comparison and the reliability of the drawn conclusions.
Datasets.
We evaluated the performance of HTMformer on eight widely used datasets, including Weather, Traffic, Electricity, ETTh2 adopted in Autoformer (Wu et al., 2021), Solar-Energy proposed in LSTNet (Lai et al., 2018), and three PEMS datasets (PEMS03, PEMS04, PEMS08) investigated in SCINet (Liu et al., 2022). These datasets span diverse domains and sampling frequencies, providing a comprehensive experimental scenario for model evaluation.
Baselines and metrics.
We employ eight carefully selected state-of-the-art Transformer-based models and widely recognized non-Transformer-based models in the time series forecasting domain as baselines, encompassing iTransformer (Liu et al., 2024b), PatchTST (Nie et al., 2023), FEDformer (Zhou et al., 2022), Dlinear (Zeng et al., 2023), WPMixer (Murad et al., 2025), MultiPatchFormer (Naghashi et al., 2025), TimeMixer (Wang et al., 2024a), and SegRNN (Lin et al., 2023). Mean Squared Error (MSE) and Mean Absolute Error (MAE) are utilized as the assessment criteria.
@@ 4.2 Performance Promotion with HTME
To assess the effectiveness and scalability of HTME, it is integrated into the Transformer (Vaswani et al., 2017) and its variants, as shown in Table 1. Specifically, HTME achieves consistent improvements across different Transformer variants, with average performance gains of 35.8%, 34.3%, 43.6%, and 31.9% for Transformer, Reformer (Kitaev et al., 2020), Informer (Zhou et al., 2021), Flowformer (Ma et al., 2023), and Flashformer (Dao et al., 2022), respectively. These results demonstrate that HTME captures more informative time series representations and significantly strengthens the modeling capacity of Transformer-based forecasters. More importantly, the temporal feature extraction module and the multivariate feature extraction module within the HTME layer operate as entirely decoupled components. Consequently, the multivariate representations produced by HTME can be seamlessly integrated with other embedding schemes that emphasize temporal characteristics, thereby further enhancing the applicability and versatility of HTME. The HTME framework is broadly applicable to Transformer-based forecasting architectures, allowing for flexible utilization of various attention mechanisms tailored to different application scenarios.
@@
Models Transformer Reformer Informer Flowformer Flashformer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Electricity Original
@@ 4.3 Long-term Forecasting
Long-term forecasting holds significant importance in various application domains, including meteorology, traffic management, and energy utilization. We conduct a comprehensive evaluation of the effectiveness of HTMformer in long-term forecasting. Table 2 presents the experimental forecasting results. Within each row, the lowest averaged MSE and MAE values across four prediction horizons are highlighted in red, and the second-lowest values are underscored in blue. HTMformer outperforms other models in long-term forecasting across various datasets, especially on high-dimensional datasets such as Traffic. We further provide a visualization of the long-term forecasting results produced by HTMformer (see Appendix I). HTMformer effectively captures the periodic characteristics of time series, delivers accurate trend forecasts for complex sequences, and demonstrates strong robustness against large-scale fluctuations. In conclusion, HTMformer demonstrates significant performance improvements and tangible practical benefits in long-term forecasting tasks.
Table 2: Average performance for long-term forecasting over prediction horizons $H\in\{96,192,336,720\}$ with fixed lookback $L=96$. Notably, MPFormer is short for MultiPatchFormer. Full results are listed in Appendix E.2.
@@ 4.4 Short-term Forecasting
Short-term time series forecasting tasks are also prevalent across a wide range of application domains. Table 3 presents the experimental results on the PEMS (PEMS03, PEMS04, and PEMS08) datasets. The proposed HTMformer demonstrates better performance across all three PEMS datasets. It outperforms the second-best model, MultiPatchFormer, yielding a 21.7% reduction in MSE and a 12.0% reduction in MAE, respectively. We attribute this performance gain to the exploitation of rich multivariate representations in HTME. Since short-term forecasting, constrained by the limited prediction length, does not entail intricate long temporal dependencies, it renders inter-variable correlations more salient. These results validate the robust performance of the HTMformer on short-term forecasting tasks, further complementing its strong performance in long-term forecasting.
@@
Compared with iTransformer (Liu et al., 2024b), which explicitly models multivariate correlations, and PatchTST (Nie et al., 2023), which excels at capturing fine-grained temporal patterns, HTMformer achieves broadly superior performance on both long-term and short-term forecasting tasks. Even when compared with the previously best-performing Transformer-based forecaster, MultiPatchFormer, HTMformer consistently achieves either superior or comparable predictive accuracy. This performance gain is particularly pronounced on high-dimensional time series datasets, such as Traffic and Solar-Energy. Such effectiveness is attributed to the use of the HTME extractor, which facilitates multi-dimensional feature fusion in time series. HTME adaptively integrates temporal and multivariate representations to achieve comprehensive, multidimensional feature complementarity. Furthermore, it encapsulates richer and more discriminative semantic representations, enabling Transformer-based forecasters to more fully leverage sequence information. This demonstrates the broad applicability across diverse datasets, thereby highlighting the effectiveness of HTME.
@@ 5.1 Ablation Studies
To evaluate the contribution and reliability of each module within HTMformer, we conducted ablation studies (see Appendix F). Specifically, we examined two different variants of HTMformer: 1) HTMformerV1, which utilizes only the temporal feature extraction module of the HTME extractor; 2) HTMformerV2, which utilizes only the multivariate feature extraction module of the HTME extractor. We also examined another variant 3) iTransformerV3 (iTransV3), which integrates the multivariate feature extraction module of the HTME extractor into the embedding layer of iTransformer. We adopt iTransformer (Liu et al., 2024b) as the state-of-the-art baseline. Sharing a highly similar architectural design with HTMformer, both models employ a dedicated embedding module and process inverted inputs before feeding them into the Transformer encoder. However, in iTransformer, the embedding layer conducts feature extraction exclusively along the temporal dimension. Notably, since all experiments were performed under identical configurations, the HTMformer variants are directly comparable to the other models. We also adopt MultiPatchFormer (Naghashi et al., 2025), attaining the second-best performance in various forecasting tasks, as a representative proxy for other forecasting models. Specific results are documented in Table 4.
@@
Compared with other time series forecasting models, HTMformerV1 already outperforms several baselines and suffers only a marginal accuracy drop even when benchmarked against state-of-the-art approaches. Blue numbers indicate the cases where HTMformerV1 outperforms iTransformer. This observation suggests that our temporal feature extraction module has effectively captured informative temporal patterns.
Effect of the multivariate features.
To facilitate an intuitive comparison, we evaluate HTMformerV1 against HTMformerV2, with numerical values underlined when HTMformerV2 achieves better performance over HTMformerV1. Across most datasets, particularly those with a large number of variables, HTMformerV2 performs substantially worse than other time series forecasting models, highlighting that effective forecasting primarily depends on accurate temporal dependency modeling. Nevertheless, on certain datasets, HTMformerV2 attains superior or comparable results relative to most baseline models. This emphasizes that multivariate features are crucial and indispensable sources of information in time series forecasting.
@@
We further conduct comparative evaluations between HTMformer and HTMformerV1, as well as between iTransformerV1 and the original iTransformer. Red numbers indicate cases where the versions augmented with multivariate integration in the embedding layer outperform their original counterparts. Across most datasets, both HTMformer and iTransformerV1 consistently achieve superior performance. This provides strong evidence for the effectiveness of our proposed strategy for incorporating multivariate features into embedding representations.
@@ 5.2 Model Efficiency
To assess the computational efficiency of HTMformer, we perform a comparative analysis of its training time, GPU memory footprint, and parameter count against state-of-the-art models, including MultiPatchFormer (Naghashi et al., 2025), WPMixer (Murad et al., 2025), and PatchTST (Nie et al., 2023). All models are evaluated on the same GPU under identical lookback windows, prediction horizons, and batch sizes to ensure fair comparison. As shown in Figure 4, using the state-of-the-art MultiPatchFormer as the reference baseline, HTMformer achieves a training runtime approximately one-third of that of MultiPatchFormer, while requiring only 20% to 45% of its GPU memory footprint. Moreover, HTMformer’s model parameters are just half those of MultiPatchFormer. Importantly, these gains are achieved without sacrificing predictive accuracy, as HTMformer consistently attains superior or comparable MSE and MAE across eight benchmark datasets. These results indicate that HTMformer enables faster training and inference under the same hardware conditions, and its smaller memory footprint also enhances the model’s applicability (see Appendix G).
@@
It is noteworthy that the ablation studies and the long-term forecasting experiments were conducted under identical configurations, rendering the results directly comparable. We observe that for time series data, using patching and convolutions for local patterns and linear layers for long-range dependencies is generally sufficient for effective temporal modeling. Excessive modeling of temporal dependencies yields limited gains in forecasting accuracy while substantially increasing compute and memory overhead. In this study, we adopt a hybrid design in the embedding layer, employing two lightweight models to extract multidimensional features. The resulting hybrid embeddings extract richer informative representations from the sequences, thereby assisting Transformer-based forecasters in modeling time series. This approach not only improves predictive performance but also markedly reduces both computational demand and memory footprint.
@@ 6 Conclusion and Limitations
Motivated by the inherent properties of multivariate time series, we introduce the HTME strategy. The HTME module jointly encodes temporal dynamics and inter-variate dependencies present in the raw sequences, yielding embedding representations that are richer and more informative than those produced by prior designs. By inverting the input representation, the attention mechanism is empowered to more effectively model inter-channel dependencies. Experimental results demonstrate that HTMformer delivers state-of-the-art forecasting performance with faster training and a reduced GPU memory footprint, thereby validating the effectiveness of injecting multivariate information into the embedding space. Moreover, the proposed HTME module in HTMformer is seamlessly compatible with diverse Transformer architectures, underscoring its strong scalability.
Through numerous experiments, this paper further highlights that incorporating both temporal and multivariate information enables Transformer-based predictors to make more accurate predictions. The detailed discussion is provided in Appendix J. However, time series data exhibit complex correlations, and simply adding multivariate features to the temporal dimension through the embedding layer cannot fully model such a complex pattern. Jointly modeling spatiotemporal dependencies in time series remains an open problem that requires further exploration. Future improvements in Transformers can focus on devising methods to efficiently capture multivariate correlations.
@@ 10 Acknowledgements
Appendix A Experiments of Different Embedding Methods
To demonstrate that Transformer-based forecasters have limited ability to understand sequences and the necessity of deeply mining sequence features in the embedding layer, we compared the prediction performance of Transformer (Vaswani et al., 2017), iTransformer (Liu et al., 2024b), and PatchTST (Nie et al., 2023) with different embedding approaches on four datasets: ECL, Weather, ETTh2, and Solar. The input sequence length is set to 48, 96, 144, 192, 240, and 288, with the prediction length being 96. The experimental results are shown in Figure 5.
Experimental results reveal that increasing the input sequence length initially leads to a significant decline in prediction loss. However, as the sequence length continues to grow, this benefit markedly lessens, and in some cases, prediction loss may even increase. These findings suggest that Transformer-based models exhibit limitations in proficiency in capturing effective features from raw series data. Simply increasing the lookback window does not consistently capture more informative patterns and will also incur additional computational overhead. This is because time series forecasting focuses on capturing short-term dependencies (Kim et al., 2025) and the temporal dimension inherently offers limited contextual information. Therefore, traditional Transformer-based forecasters encounter bottlenecks in reliably improving prediction accuracy. Notably, employing more sophisticated embedding methods to capture complex spatiotemporal dependencies can effectively address this problem. Motivated by these findings, we propose a novel embedding strategy that moderately extracts features along the temporal dimension to balance prediction accuracy and computational complexity, while integrating multivariate features to enhance informative content, thereby enabling the efficient acquisition of high-quality embeddings. Combined with input inversion, this approach brings comprehensive improvements to Transformer-based forecasters.
@@ Appendix A Experiments of Different Embedding Methods
Appendix B Limitations of Conventional Embedding
Time series data inherently exhibits both temporal and multivariate (spatial) characteristics. A growing body of research has demonstrated the effectiveness of explicitly capturing intra- and inter-channel dependencies (Zhang & Yan, 2023; Liu et al., 2024b; a). In prevailing Transformer-based forecasting frameworks, the embedding stage typically consists of three main components: timestamp embedding, positional embedding, and time series convolution. These components are primarily designed to capture temporal patterns while often overlooking the intrinsic correlations among different variables within the data. In fact, many types of time series data exhibit significant multivariate correlations, such as traffic flow and weather forecasting (Li et al., 2018; Wu et al., 2020b), where sensors are installed at fixed spatial locations. However, conventional embeddings capture only temporal patterns, while the subsequent attention layer also primarily models temporal dependencies, thereby overlooking multivariate dependencies and leading to information loss. Due to the overemphasis on temporal modeling and the loss of multivariate information under such designs, existing Transformer-based forecasters have encountered significant bottlenecks in capturing the contextual relationships essential for accurate multivariate time series forecasting.
Furthermore, the temporal feature extraction approach of the embedding layer, inspired by large language models (LLMs), is ill-suited for time series analysis. Unlike the Transformers used in natural language and vision tasks (Kirillov et al., 2023; OpenAI, 2023), which require learning dependencies among thousands to millions of tokens, Transformers in time series forecasting generally operate within relatively limited contexts, involving no more than hundreds of time series tokens. A smaller number of tokens makes it difficult for the attention mechanism to learn clear temporal correlations. However, time series convolution is limited to extracting local patterns from the original sequence and lacks the capability for fine-grained feature extraction. Additionally, PRformer (Yu et al., 2025) highlights that traditional Transformer algorithms attempt to learn sequence variation patterns by combining positional encoding with multivariate data at different time instants. However, real-world time series are frequently non-stationary, and the assumption of translation invariance makes it difficult for the Transformer’s self-attention mechanism to learn correct temporal dependencies. Our experimental results (see Appendix A) indicate that the quality of the embeddings has a substantial impact on the performance of Transformer-based forecasters. Therefore, it is imperative to redesign the embedding model to capture more effective representations.
The HTME extractor discards positional encoding and treats timestamp information as a separate variable, concatenated to the time series inputs. This method has enhanced the model’s ability to learn the relationship between the sequence and time. It employs patch-based operations and convolution to more effectively capture short-term dependencies, and a linear layer to capture long-term correlations. This method can extract temporal features more effectively and is better suited for Transformer-based architectures in time series forecasting tasks. We further employ a well-designed multivariate feature extractor to enrich the embedding representations with multivariate features. Notably, we do not explicitly model multivariate correlations, as doing so would incur substantial computational overhead. Finally, it incorporates multivariate features into temporal features to complement and adaptively balances the contributions of the two feature components, thereby enhancing the scalability of the embeddings. The resulting HTME incorporates rich temporal information and thoroughly accounts for multivariate correlation effects. It is more effective than existing embeddings in extracting informative representations from raw time series. This approach has already achieved good results in temporal feature extraction, making it unnecessary to overemphasize temporal correlations through the attention mechanism. So HTMformer utilizes inverted input to enable the attention mechanism to capture correlations among multiple features. iTransformer (Liu et al., 2024b) demonstrates that this approach leads to improved performance.
Appendix C Time and Memory Complexity
Table 5 provides a detailed comparison of the time complexity and memory usage across various Transformer-based forecasters. $L$ denotes the input sequence length, $N$ denotes the number of variables, and $S$ denotes the stride length of PatchTST.
Table 5: Time and memory complexity of different Transformer-based forecasters.
Methods Time Memory
Transformer $O(L^{2})$ $O(L^{2})$
Informer $O(L\log L)$ $O(L\log L)$
Reformer $O(L\log L)$ $O(L\log L)$
PatchTST $O({L^{2}}/{S^{2}})$ $O({L^{2}}/{S^{2}})$
iTransformer $O(N^{2})$ $O(N^{2})$
HTMformer $O(N^{2})$ $O(N^{2})$

### v3

@@ 4.1 Experiment Settings
All models are evaluated under identical settings, following the Time-Series-Library (wang2024tssurvey), to ensure a fair comparison and reliable conclusions.. We evaluated the performance of HTMformer on eight widely used datasets, including Weather, Traffic, Electricity, ETTh2 adopted in Autoformer (wu2021autoformer), Solar-Energy proposed in LSTNet (lai2018modeling), and three PEMS datasets (PEMS03, PEMS04, PEMS08) investigated in SCINet (liu2022scinet). These datasets span diverse domains and sampling frequencies, providing a comprehensive experimental scenario for model evaluation. We employ eight carefully selected state-of-the-art models from three different architectures.
@@ 4.2 Performance Promotion with HTME
To assess the effectiveness and scalability of HTME, the vanilla self-attention mechanism (vaswani2017attention) and its variants are integrtated into the HTMformer framework, as shown in Table 1. Specifically, HTMformers achieve consistent improvements across different variants, with average performance gains of 35.8%, 34.6%, 43.6%, 31.9%,and 33.1% for Transformer, Reformer (kitaev2020reformer), Informer (zhou2021informer), Flowformer (ma2023flowformer), and Flashformer (dao2022flashattention), respectively.
@@
Models Transformer Reformer Informer Flowformer Flashformer iTransformer DLinear NLinear SegRNN
ECL Original
@@
4.6%
4.2%
3.0%
1.7%
To demonstrate the strong scalability of HTME strategy, we apply it to diverse architectures, including the Transformer-based iTransformer, the linear-based DLinear, NLinear and the RNN-based SegRNN. For these models, we supplement the original input with multivariate features. The results show that HTME consistently improves the forecasting performance across all these models.
The HTME strategy consistently boosts the performance of a wide range of forecasting models. More importantly, the multivariate representations produced by HTME can be seamlessly integrated with other embedding schemes that emphasize temporal characteristics. These further enhance the effectiveness and scalability of HTME.
@@ 4.3 Long-term Forecasting
Long-term forecasting is critical for real-world applications such as meteorology, traffic management, and energy systems. Table 2 presents the experimental forecasting results. Within each row, the lowest averaged MSE and MAE values across four prediction horizons are highlighted in red, and the second-lowest values are underscored in blue. HTMformer achieves the best average performance on most datasets, with particularly strong gains on high-dimensional benchmarks such as Traffic. We further provide qualitative visualizations of the predicted trajectories (see Appendix J). Overall, HTMformer delivers significant improvements in long-term forecasting, highlighting its practical value.
Table 2: Average performance for long-term forecasting over prediction horizons $H\in\{96,192,336,720\}$. MPFormer is MultiPatchFormer. Full results are listed in Appendix E.2.
@@
nie2023time
zeng2023dlinear
zhou2022fedformer
naghashi2025multipatchformer
murad2024wp
wang2023timemixer
lin2023segrnn
@@ 4.4 Short-term Forecasting
Short-term time series forecasting tasks is also prevalent in many real-world applications. Table 3 presents the experimental results on the PEMS (PEMS03, PEMS04, and PEMS08) datasets. The proposed HTMformer demonstrates BEST performance across all three PEMS datasets, outperforming the second-best model, MultiPatchFormer, with a 21.7% reduction in MSE and a 12.0% reduction in MAE. We attribute this performance gain to the exploitation of rich multivariate representations in HTME. Since short-term forecasting, constrained by the limited prediction length, does not entail intricate long temporal dependencies, it renders inter-variable correlations more salient. These results further demonstrate HTMformer’s strong and robust performance, complementing its long-term results.
@@
nie2023time
zeng2023dlinear
zhou2022fedformer
naghashi2025multipatchformer
murad2024wp
wang2023timemixer
@@ 5.1 Ablation Studies
To evaluate the contribution of each component in HTME, we conduct ablation studies (see Appendix F). Specifically, we examined three different variants of HTMProductors: 1) Productor, the baseline forecasting model. 2) ProductorV1, which utilizes only the multivariate feature extraction module of the HTME extractor. 3) ProductorV2, which utilizes the multivariate feature extraction module of the HTME extractor to supplement the embedding representations. We further introduce two variants: HTMformer, builds upon iTransformer by incorporating the complete HTME extractor; HTMformerV1, replaces the original embedding layer in iTransformer with the temporal feature extractor from HTME. We adopt iTransformer (liu2024itransformer) as the state-of-the-art baseline. The results are summarized in Table 4.
@@
The experimental results show that HTMformer-A and iTransformer outperform iTransformerV1 in most cases. This suggests that the primary features of time series are embedded within the temporal dimension. Furthermore, HTMformer-A demonstrates superior performance over iTransformer across the majority of which showwd in blue. This indicates that deeply mining temporal patterns can further enhance prediction performance. These findings underscore the dominant role of temporal features in time series forecasting.
@@
On the PEMS dataset, which exhibits strong multivariate correlations, PredictorV2 achieved superior performance. This indicates that multivariate correlation is a significant pattern in prediction. However, for most datasets, multivariate feature extraction can introduce substantial noise.
Effect of the hybrid strategy.
We further conduct comparative evaluations between Predictors and PredictorV1. Red numbers indicate cases where the versions augmented with multivariate integration in the embedding layer outperform their original counterparts. Across most datasets, PredictorV1 consistently achieved superior performance. This provides strong evidence for the effectiveness of our proposed strategy for incorporating multivariate features into embedding representations. The multivariate features consistently enriched the temporal feature embeddings.
@@ 5.2 Model Efficiency
To assess the computational efficiency of HTME, we compare the floating-point operations (FLOPs) and the number of parameters of iTransformer, DLinear, RLinear, and SegRNN against their HTME versions. Owing to its carefully designed architecture, the computational overhead introduced by the HTME multivariate feature extraction module is nearly constant: for complex Transformer-based models it adds negligible extra cost, and for simple models the overall overhead remains modest.
We perform a comparative analysis of HTMformer against MultiPatchFormer (naghashi2025multipatchformer), and PatchTST (nie2023time). As shown in Figure 5, taking the state-of-the-art MultiPatchFormer as the reference baseline, HTMformer reduces training runtime to roughly one third while using only 20%–45% of the GPU memory. Moreover, HTMformer’s model parameters are just half those of MultiPatchFormer. Importantly, these gains are achieved without sacrificing predictive accuracy, as HTMformer consistently attains superior or comparable MSE and MAE across eight benchmark datasets. The HTME strategy enables simple backbones to achieve performance comparable to much more complex architectures by introducing only a lightweight module (see Appendix G).
@@
5.3 Analysis of HTME
To better understand the feature space induced by HTME, we visualize the parameters defined in Equation (6), which maps multivariate feature embeddings onto the temporal feature space. The visualization shows that temporal features are consistently assigned substantially larger weights than multivariate features. The vast majority of features in time series are contained within the temporal dimension (hyndman2021forecasting). The multivariate dimension also contains a number of features (Pravilovic2013SpatialForecasting). However, it does not warrant the use of an additional complex module for its modeling. In the visualization, temporal features exhibit a line-shaped distribution, indicating a consistent temporal correlation within the series. In contrast, multivariate features appear as scattered point-like clusters, emphasizing a few salient variables. With this design, HTME augments temporal embeddings with multivariate signals: the embeddings fed into the main backbone are still treated as temporal features, but they are now enriched with additional semantic information, thereby enhancing the backbone’s modeling capacity. This approach avoids introducing extra modules into the backbone network and ensures the consistency of the feature space (see Appendix H).
@@ 6 Conclusion and Limitations
In this work, we propose the HTME strategy. The HTME module jointly encodes temporal dynamics and inter-variate dependencies present in the raw sequences, yielding embedding representations that are richer and more informative than those produced by prior designs. This allows the predictor to avoid introducing additional modules for modeling multivariate correlations. Consequently, HTME consistently enhances the modeling capabilities of various models, enabling lightweight models to achieve performance comparable to that of complex ones.
However, time series data exhibit complex correlations, and simply adding multivariate features to the temporal dimension through the embedding layer cannot fully model such a complex pattern. Jointly modeling spatiotemporal dependencies in time series remains an open problem that requires further exploration.
@@ 10 Acknowledgements
The appendix is divided into several sections, each giving extra information and details.
Appendix A Discussion on Multivariate Feature Extraction
Recent studies (han2023capacity; li2023revisiting) have found that Channel Dependence ideally gains from higher capacity, while Channel Independence can significantly enhance performance due to sample scarcity. As most current forecasting benchmarks are not sufficiently large, the complex design of the Channel Dependence model makes it prone to capturing spurious correlations from noise and overemphasizing coincidental patterns in the training data. In time series forecasting, an increasing number of models employ Channel Independence (nie2023time) or exclusively focus on modeling the temporal dimension. These models achieved good performance. However, they inevitably overlook the multivariate correlations, resulting in the loss of some information. We believe that multivariate correlations should not be ignored. Past hybrid models explicitly model multivariate correlations by incorporating various extra and complex modules within their backbone networks. This strategy mixes temporal and multivariate features within the same space. However, constrained by the sample size, the multivariate extraction process can introduce significant noise and spurious patterns. Moreover, the additional multivariate features dilute the temporal features that should be dominant, leading to a disorganized feature space. Moreover, the additional and complex module for multivariate feature extraction introduces significant computational overhead.
Appendix B Experiments of Different Embedding Methods
To demonstrate that Transformer-based forecasters have limited ability to understand sequences and the necessity of deeply mining sequence features in the embedding layer, we compared the prediction performance of Transformer (vaswani2017attention), iTransformer (liu2024itransformer), and PatchTST (nie2023time) with different embedding approaches on four datasets: ECL, Weather, ETTh2, and Solar. The input sequence length is set to 48, 96, 144, 192, 240, and 288, with the prediction length being 96. The experimental results are shown in Figure 6.
Experimental results reveal that employing more sophisticated embedding methods to capture complex spatiotemporal dependencies can effectively enhance the performance of Transformers. These findings suggest that Transformer-based models exhibit limitations in proficiency in capturing effective features. This limitation arises because the embedding layer fails to effectively represent the complex characteristics of time series. Constructing a sophisticated embedding method to extract semantically richer representations from the raw sequence can significantly enhance the performance of Transformers.
We also noticed that increasing the input sequence length initially leads to a significant decline in prediction loss. However, as the sequence length continues to grow, this benefit markedly lessens, and in some cases, prediction loss may even increase. Simply increasing the lookback window does not consistently capture more informative patterns and will also incur additional computational overhead. This is because time series forecasting focuses on capturing short-term dependencies (kim2025comprehensivesurveydeeplearning) and the temporal dimension inherently offers limited contextual information. Therefore, we argue that it is highly beneficial to consider multivariate correlations during the embedding stage. Doing so can greatly enrich the information content of the embedded representations and, in turn, enhance the modeling capability of subsequent modules.
Motivated by these findings, we propose a novel strategy that moderately extracts features along the multivariate dimension to enhance informative content, thereby enabling the efficient acquisition of high-quality embeddings.
@@ Appendix B Experiments of Different Embedding Methods
Appendix C Detail of HTME
To demonstrate the correctness of the HTME extractor, we provide a detailed theoretical explanation for each of its steps. The HTME temporal feature extraction module consists of four steps: Patching, Stacking, short-term feature extraction, and long-term feature extraction. The purpose of the Patching step is to segment the input into short-term temporal sequences. The Stacking step is inspired by the concept of channel independence, which has been proven to be more effective (see Appendix A). Time series data inherently exhibit strong short-term correlations and weaker long-term dependencies (see Appendix C). Motivated by this, we first employ convolutional layers to extract multi-scale representations from the short-term sequences, followed by a linear layer to capture local features. Subsequently, another linear layer is used to model the long-term dependencies among these features. This hierarchical process yields a more informative temporal feature representation.
The key to multivariate feature extraction is to reduce overfitting and mitigate noise (see Appendix A). For the multivariate feature extraction module, we adopt the exact same logic and methodology. It is worth noting that we cannot simply treat the variables as being independent of the temporal dimension. We concatenate the multivariate data over short time windows and then use a linear layer to extract features, to account for the effect of time delays. A change occurring at one node may exert influence on other nodes at the subsequent time step. This extraction strategy, which fuses short timesteps, enables the model to deeply consider the complex multivariate correlations within time series, thereby effectively suppressing over-attention to coincidental patterns.
$N^{(L-K),128}$ $=\text{Linear}(T^{(L-K),K\times N})$
Subsequently, we employ a limited number of convolutional layers to extract multivariate features at different scales. Fewer convolutional kernels allows the model to focus on significant patterns, thereby substantially mitigating overfitting. Then, a specialized linear layer that GRU network integrated multi-scale features to further extract effective features.
$B^{N}$ $=\text{GRU}(\text{Conv}(N^{(L-K),128}))$
Finally, small convolutional kernels are used to fine-tune the multivariate features along the temporal dimension, which prevents the model from learning fixed patterns.
$V_{\text{out}}$ $=\text{Conv}(B^{N})$
Finally, we employ a parameter to adjust the weights of the two spaces. This parameter is initialized with a large value to guide subsequent modules to focus primarily on temporal features. During training, HTME adaptively considers the influence of multivariate features to generate high-quality embedding representations.
$Y_{\text{in}}$ $=\alpha D_{\text{out}}+(1-\alpha)V_{\text{out}},$